OpenStudy (anonymous):

Movies often use models to create special effects. In the movie Son of Godzilla, Godzilla appears to be 67 feet tall and was played by a man in a 7.5 ft rubber suit. Godzilla’s son appeared to be 45 feet tall in the movie; how tall was the rubber suit for the actor playing the son? (Round to the nearest hundredth)

HanAkoSolo (jamierox4ev3r):

in order to solve this, you would create ratios between the two situations

HanAkoSolo (jamierox4ev3r):

\(\frac{67}{7.5}\)=\(\frac{45}{x}\)

HanAkoSolo (jamierox4ev3r):

idk if you can read that, but is 67/7.5=45/x

OpenStudy (anonymous):

ok.. im with yah

HanAkoSolo (jamierox4ev3r):

then, cross multply between the two

HanAkoSolo (jamierox4ev3r):

that gives you 67x=337.5

HanAkoSolo (jamierox4ev3r):

then divide 337.5 by 67, and you get 5.0373131433....... you can round that to the nearest foot, and you get 5 ft.

OpenStudy (anonymous):

ok. but it just says round to the nearest hundredth. that would be 5.03 right?

HanAkoSolo (jamierox4ev3r):

no, that would be 5.04, because you have to take into account the 7 next to the three. Good job, i think you have a good idea on how to solve this problem :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):



HanAkoSolo (jamierox4ev3r):

haha \(\huge\ddot\smile\)